OpenStudy (anonymous):

please help i don't understand and I'm gonna fail!! Task 1 Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f x x a b g x c x d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

the equations are f(x)=x+a/b and g(x)=cx-d

OpenStudy (anonymous):

@Maria977

OpenStudy (anonymous):

@Arielle

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

any numbers?

OpenStudy (anonymous):

Hello

OpenStudy (misty1212):

lets start with \[g(x)=cx-d\] and plug in some numbers here

OpenStudy (misty1212):

lets not make it too hard \[g(x)=3x-4\] now lets find \(g^{-1}(x)\) by solving \[x=3y-4\] for \(y\)

OpenStudy (misty1212):

add \(4\) get \(x+4=3y\) then divide by \(3\) to get \[f(x)=\frac{x+4}{3}\]

OpenStudy (misty1212):

so you can say "\(a=4,b=3,c=4,d=3\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the equations would be f(x)=x+4divided by 3 and g(x)=4(x)-3

OpenStudy (anonymous):

@misty1212